MULTI-RESOLUTION STUDIES OF COMPLEX MOLECULES IN HIGH MASS STAR FORMING REGIONS

\underline{DOUGLAS N. FRIEDEL}, Department of Astronomy, University of Illinois, 1002 W. Green St., Urbana, IL 61801.

After our successful molecular studies of the high mass star forming region Orion-KL (Friedel & Widicus Weaver 2012, ApJS, 201, 17; Widicus Weaver & Friedel 2012, ApJS, 201, 16) we have observed three more high mass star forming regions (W3, W75N, and W51). The purpose of these observations is to further constrain the physical conditions under which the different complex organic molecules form and survive. The results of these observations as well as their impact on current molecular formation models will be presented.